### v1

# Performance of multiple filter-cavity schemes for frequency-dependent squeezing in gravitational-wave detectors

Abstract
Gravitational-wave detectors use state-of-the-art quantum technologies to circumvent vacuum fluctuations via squeezed states of light. Future detectors such as Einstein Telescope may require the use of two filter cavities or a 3-mirror, coupled filter cavity to achieve a complex rotation of the squeezing ellipse in order to reduce the quantum noise over the whole detector bandwidth. In this work, we compare the theoretical feasibility and performances of these two systems and their resilience with respect to different degradation sources (optical losses, mismatching, locking precision). We provide both analytical models and numerical insights. We extend previous analysis on squeezing degradation and find that the coupled cavity scheme provides similar or better performances than the two-cavity option, in terms of resilience with respect to imperfections and optical losses. We propose a possible two-step implementation scheme for Einstein Telescope using a single filter cavity that can be possibly upgraded into a coupled filter cavity.
preprint: APS/123-QED
I Introduction
Over the past decade, LIGO, Virgo and KAGRA have opened the era of gravitational-wave (GW) astronomy, detecting hundreds of signals from merging compact binary objects, including black-hole binaries, neutron-star binaries and black-hole-neutron-star systems [1]. These observations have led to a wide range of scientific insights in fields as diverse as general relativity, astrophysics and cosmology [2].
Quantum noise limits the ultimate sensitivity of GW detectors. Such noise is due to vacuum fluctuations of the electromagnetic field entering the detector through the output port. Quantum noise can be mitigated by generating and injecting squeezed vacuum through this port [3, 4]. Because of the inherent frequency response of the interferometer, and the ponderomotive coupling with the mirrors, such a squeezed vacuum state needs to acquire a frequency dependence in order to optimally reduce quantum noise in the whole detector bandwidth. Precisely, the interferometer rotates the squeezed quadratures in a frequency-dependent manner. Fig. 1 shows the simplified optical scheme of such a quantum-enhanced GW interferometer.
Figure 1: Simplified optical scheme for dual-recycled GW interferometer with Frequency-Dependent Squeezing injected through the output port. Current and future generations of GW detectors use the same simplified scheme, but ET-LF operates in a detuned configuration.
Frequency-Dependent Squeezing (FDS) is achieved by first generating a frequency-independent squeezed vacuum state, then passing it through one or several filter cavities to rotate its quadratures, thus compensating the response of the interferometer [5, 6, 7]. Recently, both Advanced Virgo and Advanced LIGO demonstrated FDS using a single filter cavity on site [8, 9], because the interferometers are operated in a broadband configuration [10]. In particular, the implementation of FDS in Advanced LIGO improved the detection rate up to 65% [11] and allowed operation below the standard quantum limit [12]. While not currently implemented in LIGO or Virgo, note that an alternative scheme – called EPR squeezing – which does not directly require a filter cavity to achieve FDS, has been theoretically proposed [13] and demonstrated in proof-of-principle experiments [14, 15].
At present, “third-generation” (3G) detectors, to be built in new infrastructures, are being studied and aim to increase sensitivity by a factor of $\sim 10$ compared to existing detectors. They will expand the science cases of GW, addressing fundamental open questions such as the nature of gravity and dark energy, the properties of nuclear matter, and the formation of neutron stars and black holes throughout cosmic history. Einstein Telescope, one of such 3G detectors, is composed of two interferometer designs: a “hot” detector, optimized for high-frequency sensitivity, and a “cold” detector (ET-LF), operating at cryogenic temperatures for improved low-frequency sensitivity [16]. This work focuses on the quantum noise reduction (QNR) in ET-LF as a 10-km, single L-shaped (right angle between the arms), Michelson interferometer.
While ET-LF’s sensitivity will be limited by quantum noise, similar to current GW detectors, its mitigation will require two filter cavities (referenced as 2FC in the following) [17]. This is because ET-LF will operate in a detuned (non-broadband) configuration [18, 19]. Among alternatives schemes that do not require two filter cavities, such as quantum teleportation [20], we consider here a three-mirror cavity, otherwise known as a coupled filter cavity (henceforth referenced as CFC), which has been first proposed in [21]. A comparison of 2FC vs. CFC in the lossless case was performed in [22], where it was identified that the equivalence between the two schemes imposed a value for the CFC middle mirror transmissivity apparently smaller than current state of the art techniques in mirror coating.
In this paper, we review the equivalency between 2FC and CFC, and address the issue of the middle mirror transmission through compensating mechanisms (Section II). In particular, we show that, for longer filter cavities than that studied in [22], its value is actually readily achievable. We then perform a complete comparison between 2FC and CFC configuration in terms of squeezing degradation budget, in particular including the effects of optical loss, mismatch and phase noise (Section III). It is shown that, theoretically, CFC may provide better resilience to mode-matching and phase noise than 2FC. Finally, we introduce a tuned configuration of ET-LF (with single filter cavity) as an intermediary step before achieving detuned ET-LF, and study its astrophysical merit (Section IV).
II Proposed configurations
We review below the two optical configurations for producing FDS which are the subject of this paper. Let $L_{\text{tot}}$ be the total length of the filter cavities. They are shown in Fig. 2.
  • 2FC configuration with cavity lengths $L_{1}$, $L_{2}$ such that $L_{1}+L_{2}=L_{\text{tot}}$.
  • CFC configuration with cavity lengths $L_{c}$, $L_{c}$ such that $L_{a}+L_{c}=L_{\text{tot}}$; the first sub-cavity is called ”c-cavity” and the second is the ”a-cavity”.
Figure 2: Optical schemes of the two-filter cavity (2FC, top) and coupled filter cavity (CFC, bottom) configurations. Note that, aside from the different cavity configuration, the 2FC differs from the CFC insofar at is requires mode matching lenses between the first filter cavity (FC1) and the second one (FC2), as well as an additional Faraday Isolator (FI).
Note that the main difference between these two configurations, apart from the cavities, is the additional Faraday Isolator (FI) as well as free-space mode matching elements between the first and second cavities in 2FC. These elements will further degrade squeezing compared to CFC as we will show in Section III.
II.1 Scaling law between 2FC and CFC
In the lossless case and assuming that all the cavities considered have high finesses, it is possible to obtain closed form equations relating the linewidths and the detunings of 2FC to those of CFC. They are given in [22] and recalled here:
$\gamma_{c}$ $=\gamma_{1}+\gamma_{2}$ (1)
$\gamma_{a}$ $=\frac{\gamma_{1}\gamma_{2}}{2\nu_{c}}\left[1+\left(\frac{\Delta\omega_{1}-\Delta\omega_{2}}{\gamma_{1}+\gamma_{2}}\right)^{2}\right]$ (2)
$\Delta\omega_{a}$ $=\frac{\Delta\omega_{1}\gamma_{2}+\Delta\omega_{2}\gamma_{1}}{\gamma_{1}+\gamma_{2}}$ (3)
$\Delta\omega_{c}$ $=\frac{\Delta\omega_{1}\gamma_{1}+\Delta\omega_{2}\gamma_{2}}{\gamma_{1}+\gamma_{2}}$ (4)
with $\nu_{c}=\frac{c}{2L_{c}}$ the free spectral range (FSR) of the c-cavity, $\gamma_{i}=\frac{cT_{i}}{4L_{i}}$ and $\Delta\omega_{i}$ the linewidth and detuning of cavity labeled $i$.
We recall that the values for $\gamma_{1},\gamma_{2},\Delta\omega_{1}$ and $\Delta\omega_{2}$ are fixed for a given configuration of the interferometer [17], in the lossless case. This means that a change in the cavity lengths $L_{a},L_{c}$ of the 2FC setup should be accompanied by an equal change in the input mirrors’ transmissions $T_{1},T_{2}$ to keep $\gamma_{1},\gamma_{2}$ constant. From Eq. 1, this translates into $T_{c}\propto L_{c}$ while Eq. 2 implies that $T_{a}\propto L_{a}L_{c}\propto L_{c}^{2}$, where that last scaling law reasonably assumes that $L_{a}\propto L_{c}$.
Table 1 provides the parameters which will be used in the rest of this paper (taken from [23]). We briefly comment on the chosen value of 30 ppm for round trip loss $\Lambda$. They have been estimated according to the empirical scaling law:
Λ=30 $\Lambda=30\text{ ppm}\times\left(\frac{L}{300\text{ m}}\right)^{0.3}\left(\frac{1064\text{ nm}}{\lambda}\right)^{2}$ (5)
where the values of 30 ppm has been found by simulations in [24] for a 300-m cavity illuminated with a 1064 nm laser. For the scaling low with the length of the cavity we use the result of [25]. The scaling with the laser wavelength is given by the physics of the scattering [26]. This can be considered a conservative estimation, since we expect future advances in the mirror and coating techniques.
Parameter Physical meaning Value
$\lambda$ Laser wavelength 1550 nm
$I_{\text{arm}}$ Power in the arms 18 kW
$T_{\text{arm}}$ ITM transmission 0.007
$L_{\text{arm}}$ Interferometer arm length 10 km
$m_{\text{ETM}}$ Mass of test-mass 211.3 kg
$T_{\text{SRM}}$ SRM transmission 0.2
$L_{\text{SEC}}$ SEC length 100 m
$\phi_{\text{SEC}}$ SEC detuning $\frac{\pi-0.75}{2}$ rad
$\theta_{\text{HD}}$ Homodyne angle $-0.27$ rad
$\delta\theta_{\text{HD}}$ HD RMS phase noise 10 mrad (typ.)
$\gamma_{1}$ 1st FC linewidth $2\pi\times(4.26)$ rad/s
$\gamma_{2}$ 2nd FC linewidth $2\pi\times(1.65)$ rad/s
$\Delta\omega_{1}$ 1st FC detuning $2\pi\times 19.51$ rad/s
$\Delta\omega_{2}$ 2nd FC detuning $2\pi\times(-7.65)$ rad/s
$\Lambda$ Round Trip Loss 30 ppm (typ.)
$\delta L$ FC RMS length noise 1 pm (typ.)
$\Upsilon_{I}$ Input mode mismatch 4%
$\Upsilon_{O}$ Output mode mismatch 3%
$\Upsilon_{12}$ Mode mismatch FC1-FC2 1%
$\Upsilon_{a}$ Internal mismatch (CFC) $\sim 0$%
$\Lambda_{FI}$ FI loss (double pass) 1%
$L_{\text{FC1,2}}$ Length of each FC $5$ km
$L_{\text{CFC}}$ Total length of CFC $10$ km
$T_{\text{1}}$ 1st FC input transmission $6.9\cdot 10^{-4}$
$T_{\text{2}}$ 2nd FC input transmission $1.8\cdot 10^{-3}$
$T_{\text{a}}$ CFC middle transmission $6.75\cdot 10^{-6}$
$T_{\text{c}}$ CFC input transmission $2.47\cdot 10^{-3}$
$L_{\text{FC1,2}}$ Length of each FC $1$ km
$L_{\text{CFC}}$ Total length of CFC $2$ km
$T_{\text{1}}$ 1st FC input transmission $1.4\cdot 10^{-4}$
$T_{\text{2}}$ 2nd FC input transmission $3.6\cdot 10^{-4}$
$T_{\text{a}}$ CFC middle transmission $2.7\cdot 10^{-7}$
$T_{\text{c}}$ CFC input transmission $4.95\cdot 10^{-4}$
Table 1: Parameters used in this study. For comparison, we show the parameters of 2FC and CFC if the total length is either 2 or 10 km.
II.2 The middle mirror in CFC
The middle mirror is a critical component of the CFC configuration. We first comment on concerns raised by the unusually small value for $T_{a}$. First, note that while its value in the $L_{\text{CFC}}=2$ km case ($T_{a}<0.3$ ppm) is indeed much lower than what is implemented in the current generation of GW detectors, the $L_{\text{CFC}}=10$ km case ($T_{a}=6.75$ ppm) is comparable to the transmission of the end mirror of the currently installed filter cavity in Virgo (see [27] page 350), thus technically already achieved for comparable wavelengths. Therefore, in the rest of this paper, unless explicitly specified otherwise, we will work with $L_{\text{CFC}}=10$ km.
Interestingly, if one considers the a-cavity as a single cavity, – that is, the input mirror is $T_{a}$ – then it may seem that such a cavity is loss-dominated (Λ=30 $\Lambda=30\text{ ppm}\gg T_{a}$), and that the finesse would be of the order of $10^{5}$. However, since the a-cavity is embedded in a coupled cavity system, the correct qualitative picture is to consider that the first two mirrors ($T_{c}$ and $T_{a}$) form a cavity, which is equivalent to an effective mirror with variable transmission $T_{\text{eff}}$ given as a function of the sideband frequency $\Omega$ by
$T_{\text{eff}}(\Omega)=\frac{T_{a}T_{c}(1-\Lambda)}{1+r^{2}-2r\,\cos\left(\frac{2L(\Omega-\Delta\omega_{c})}{c}\right)}$ (6)
[…]
Small deviations from the optimal middle mirror transmission $T_{\text{a}}$ as computed in Table 1 can lead to important reduction of the detector’s sensitivity. Such reduction can be partially recovered by tuning other parameters of the coupled filter cavity. However, in a realistic scenario, after the construction of the coupled filter cavity, most of the parameters cannot be adjusted. Thus, deviations in $T_{\text{a}}$ can be partially compensated by tuning only three parameters: the input squeezing angle into the CFC cavities, and the two detunings $\Delta\omega_{a}$ and $\Delta\omega_{c}$ of the sub-cavities [28].
Fig. 3 compares quantum noise reduction for 2FC with the one in the CFC case with a deviation in $T_{\text{a}}$ of 10% and 20% from its nominal value, after partial compensation. We considered all loss sources of Table 1 and we have taken the total cavity length to be $10$ km.
Therefore, by going from 20% deviation, which is the currently achievable tolerance on such a low absolute value of $T_{a}$ [29], to 10%, we can gain up to 2 dB of quantum enhancement above 20 Hz. This underlines the importance of reaching more stringent tolerances on $T_{\text{a}}$. Note that the two planar surfaces (Anti-Reflective and Highly-Reflective) of the middle mirror can be made parallel to each other to form an etalon [30]. Then, the transmissivity can be fine-tuned by controlling the temperature of the mirror, with a precision [31] that should help mitigate imperfections in the initial value of $T_{a}$.
Figure 3: Squeezing degradation for the CFC caused by a 10% and 20% deviation of the middle mirror transmission $T_{\text{a}}$ compared to the CFC model with the optimal $T_{\text{a}}$ and the two FC case. The 10% and 20% $T_{\text{a}}$ deviations are being partly compensated by adjusting the detunings of the CFC sub cavities and the injected squeezing angle. Top: Noise spectral density comparison. Bottom: Quantum enhancement in dB. Losses are being considered as detailed in Table 1.
III Comparison of the effect of different squeezing degradation sources
In this section, we provide separate comparisons of the CFC and 2FC systems to understand the contributions of the different degradation mechanisms (optical loss $\Lambda$, cavity locking precision $\delta L$, mode mismatch $(\Upsilon,\psi_{\text{mm}})$). We start with a default model for both 2FC and CFC which corresponds to their lossless configurations, and study each one of these three degradation mechanisms separately.
The general form of the quantum noise enhancement (input squeezing parameter $r$) measured at the output of a lossy passive system has the form [32]:
$\begin{split}\bar{S}[\Omega]&=\eta[\Omega]\{[(1-\Xi[\Omega])e^{-2r}+\Xi[\Omega]e^{2r}]\cos^{2}(\Delta\theta_{D}[\Omega])\\ &\quad\qquad+[(1-\Xi[\Omega])e^{2r}+\Xi[\Omega]e^{-2r}]\sin^{2}(\Delta\theta_{D}[\Omega])\}\\ &\quad+1-\eta[\Omega]\end{split}$ (7)
where $\eta[\Omega]$, $\Xi[\Omega]$ and $\Delta\theta_{D}[\Omega]$ are the three frequency-dependent figures of merit that characterize the degradation, namely the efficiency, the dephasing and the misphasing. Their mathematical definitions are provided in Appendix A. Intuitively,
  • $\eta[\Omega]$ characterizes how much input squeezing is lost to vacuum;
  • $\Xi[\Omega]$, akin to phase noise, characterizes how much sideband imbalance is induced by the system, thus how much antisqueezing is effectively coupled to squeezing through a phenomenon known as ”complex squeezing” [33];
  • $\Delta\theta_{D}[\Omega]$ encapsulates the difference of the rotation of the squeezing ellipse between the realistic and the perfect (lossless) system.
III.1 Optical Losses
Figure 4: Squeezing degradation for Round Trip Loss $\Lambda$ for 2FC (solid lines) and CFC (dashed lines), for a total length $L_{1}+L_{2}=L_{a}+L_{c}=5\text{km}+5\text{km}=10$ km. From top to bottom: Quantum noise reduction (for 10 dB of input squeezing); Relative quantum noise reduction between CFC and 2FC (a positive value means that CFC performs better than 2FC); the squeezing efficiency $\eta$, the dephasing $\sqrt{\Xi}$ and the misphasing $\Delta\theta_{D}$.
Fig. 4 compares the squeezing degradation when Round Trip Losses $\Lambda$ are added in the CFC and 2FC models (which can be physically interpreted as absorption loss on each mirror), for values of $\Lambda$ equal for all mirrors and ranging from 0 to 30 ppm.
Remarkably, we see that the degradation curves between CFC and 2FC are essentially identical for a given $\Lambda$. This was also noted in [21] but no proof was given. This is proved in Appendix B, where we we show that if CFC and 2FC have the same round trip losses (RTL), then they have the same response at the lowest order. We also see that the squeezing efficiency and dephasing reach extrema at the resonance frequencies $|\Delta\omega_{1,2}|$, while misphasing is comparatively negligible ($|\Delta\theta_{D}|\ll\sqrt{\Xi}$).
Finally, let us note that, because the middle mirror in CFC is being traversed by the optical field going from the $c-$cavity to the $a-$cavity, substrate loss may further degrade squeezing. It turns out (Appendix C) that, for current substrate loss values, this extra degradation is always inferior to 0.5 dB of squeezing thus is not significantly worsening the performance of CFC.
III.2 Mode mismatch
Mode mismatching between cavities and free-space beams has been hypothesized as the main contributor of unknown loss in current gravitational wave detectors. Its modeling is in general more involved than loss as the power of the incident beam is unitarily redistributed among higher order modes [32] and thus requires several parameters to characterize which depend on the overlap between the incident and the cavity modes.
Importantly, mode coupling can manifest in two distinct effective ways: so-called quadratic mismatch, due to waist position or size mismatch; or involve higher order aberrations, typically due to astigmatism of the optics, see [34]. The former is modeled using a unitary basis change $\bm{U}$, while the latter usually corresponds to modes which are stopped by baffles and thus can simply be modeled using an absorption coefficient $\Lambda$. In the following, we will focus our discussion on the quadratic type, that is, we assume that input, intra-cavity, and output mismatches are all modeled by unitary matrices $\bm{U}$.
We make the general assumption that mode mismatch occurs when a beam goes from free space into a cavity and vice-versa, because of limited actuation in the mode-matching telescopes and in the possible low-order aberrations that they induce. However, for coupled cavities, the mode mismatch between the two subcavities only depends on the design of the cavity mirrors, which can be made with much lower imperfections and thus this mode mismatch can made negligible (see Appendix D for a quantitative derivation).
We work in several eigenbases when propagating the modes throughout the cavities. We consider the transverse electromagnetic bases of the input field (index $I$), the eigenbasis of the first cavity (index $1$ for 2FC, $c$ for CFC), the eigenbasis of the second cavity (index $2$ for 2FC, $a$ for CFC), the eigenbasis of the output field (index $O$). In order to have tractable insights in our model, we only consider a single higher order mode that encapsulates all the mode mismatch for each basis. We regroup these two modes in a vector $\bm{a}[\Omega]=\begin{bmatrix}a_{1}[\Omega]\\ a_{2}[\Omega]\end{bmatrix}$ where the first element is the fundamental gaussian mode, while the second is the higher order mode, evaluated at the sideband frequency $\Omega$ from the carrier (see Appendix A for details). Mode mismatch is characterized by two parameters, $\Upsilon\in[0,1]$ the amplitude of the mismatch ($\Upsilon=0$ corresponds to no mismatch), and $\psi_{\text{mm}}$ the phase of the mode mismatch. For example, if we are interested in the matching between a free-space beam and a cavity eigenmode, $\psi_{\text{mm}}=0$ means that all the mismatch comes from beam waist size, while $\psi_{\text{mm}}=\pi/2$ means that all the mismatch comes from beam waist longitudinal position [35]. An intermediary value of $\psi_{\text{mm}}$ corresponds to a linear combination of waist size and position mismatch.
Figure 5: Squeezing degradation from quadratic mode mismatch in 2FC (blue, solid envelope) and CFC (red, dashed envelope), for a total length $L_{1}+L_{2}=L_{a}+L_{c}=5\text{km}+5\text{km}=10$ km. The blue and orange regions are determined by calculating squeezing curves for the values of $\Upsilon_{I}$, $\Upsilon_{O}$, $\Upsilon_{12}$ and $\Upsilon_{c}$ from Table 1, and sweeping the mode mismatch phases ($\psi_{mmO}-\psi_{mmI}$) and ($\psi_{mm12}-\psi_{mmI}$). The inter-cavity mode mismatch 2FC is $\Upsilon_{12}=0.01$. Some examples of CFC squeezing curves are plotted in orange. From top to bottom: quantum noise reduction with mismatch; ratio of the lower envelopes of CFC to 2FC; efficiency; dephasing; misphasing.
Under the quadratic mismatch model, we see in Fig. 5 that squeezing degradation presents a wideband feature characterized by the value of QNR at $\Omega\to+\infty$, and, in the worst case (lower envelopes), peaks of degradation close to the detuning frequencies $\Delta\omega_{1,2}$. To interpret these features, let us compute approximate forms for the squeezing efficiency $\eta$ and dephasing $\Xi$ at these frequencies.
We see from the numerical values of Table 1 that $\Delta\omega_{1}^{2}\gg\gamma_{1}^{2}$ and similarly $\Delta\omega_{2}^{2}\gg\gamma_{2}^{2}$, while $\Delta\omega_{1}-\Delta\omega_{2}>\gamma_{1},\gamma_{2}$. This means that we can reasonably treat both cavities separately. When the higher order mode is not resonant inside of a single cavity, its response to the squeezed mode is shown to be [32]
$F_{1FC}[\Omega]=\sqrt{1-\Upsilon_{O}}\frac{\mathrm{i}(\Omega+\Delta\omega)-\alpha\gamma}{\mathrm{i}(\Omega+\Delta\omega)+\gamma}$ (8)
where $\gamma,\Delta\omega$ are the bandwidth and detuning of the cavity, $\alpha=1-2\Upsilon_{I}+2\beta\sqrt{\Upsilon_{I}\Upsilon_{O}}\,e^{\mathrm{i}\psi_{R}}$ with $\beta=\sqrt{\frac{1-\Upsilon_{I}}{1-\Upsilon_{O}}}\simeq 1$ and $\psi_{R}=\psi_{O}+\psi_{G}-\psi_{I}$. The figures of merit are (see Eqs. 36 and 37)
$\eta[\Omega]$ $=\frac{|F_{1FC}[\Omega]|^{2}+|F_{1FC}[-\Omega]|^{2}}{2}$ (9)
$\Xi[\Omega]$ $=\frac{\left(|F_{1FC}[\Omega]|-|F_{1FC}[-\Omega]|\right)^{2}}{4\eta[\Omega]}$ (10)
At $\Omega\to\infty$, we find
$\eta[\Omega\to\infty]$ $\simeq 1-\Upsilon_{O}$ (11)
$\Xi[\Omega\to\infty]$ $=0$ (12)
which means that for CFC, which can be mapped to two single filter cavities as seen above, the high-frequency degradation is entirely determined by the output mismatch $\Upsilon_{O}$ and only shows up as loss. Numerically, for 10 dB of injected squeezing, this would correspond to a QNR of $\sim 9$ dB, consistent with the value simulated numerically. At the frequencies of the detunings, we have
$F_{1FC}[-\Delta\omega_{1,2}]$ $=-\alpha\sqrt{1-\Upsilon_{O}}$ (13)
$F_{1FC}[\Delta\omega_{1,2}]$ $=\sqrt{1-\Upsilon_{O}}.$ (14)
Thus
$\eta[\Delta\omega_{1,2}]$ $=(1-\Upsilon_{O})\frac{1+|\alpha|^{2}}{2}$ (15)
$\Xi[\Delta\omega_{1,2}]$ $=\frac{1}{2}-\frac{|\alpha|}{1+|\alpha|^{2}}$ (16)
while the misphasing is given by $\Delta\theta[-\Delta\omega_{1,2}]=\arg(\alpha)$. The lowest efficiency and highest dephasing are obtained when $\psi_{R}=\pi$, that is, the higher order mode constructively interferes to degrade squeezing. Using the values in Table 1, we estimate a QNR of $\sim 5$ dB, consistent with the plotted minima at around $\Delta\omega_{1,2}$. Thus, even for relatively low mismatch amplitude, squeezing is significantly degraded.
For 2FC, because the internal mode mismatch $\Upsilon_{12}$ is not zero in general, the expression of the response is more involved. The matrix form is:
$\begin{split}\bm{F}_{2FC}&=\bm{U}(\Upsilon_{O},\psi_{mmO})\bm{U}(\Upsilon_{I},\psi_{mmI})^{\dagger}\,\bm{U}(\Upsilon_{12},\psi_{mm12})^{\dagger}\\ &\qquad\times\bm{F}_{2}\,\bm{U}(\Upsilon_{12},\psi_{mm12})\bm{F}_{1}\bm{U}(\Upsilon_{I},\psi_{mmI})\end{split}$ (17)
where $\bm{F}_{1}=\begin{bmatrix}r_{1}&0\\ 0&1\end{bmatrix}$ with $r_{1}=-\frac{\gamma_{1}-\mathrm{i}(\Omega+\Delta\omega_{1})}{\gamma_{1}+\mathrm{i}(\Omega+\Delta\omega_{1})}$, and analogously for $\bm{F}_{2}$. We are only interested in the upper left matrix element of $\bm{F}_{2FC}$; however its general form is cumbersome. We can nevertheless determine the squeezing degradation at high frequencies $\Omega\gg\Delta\omega_{1},\Delta\omega_{2}$ as the transfer matrices simplify to $\bm{F}_{1}=\bm{F}_{2}=\bm{1}$ so $F^{(11)}_{2FC}[\pm\Omega]\underset{\Omega\to\infty}{=}\sqrt{\Upsilon_{O}}$, similarly to the CFC behavior.
To study the degradation in 2FC close to the detuning frequencies, notice in the formula Eq. 17, the first cavity “sees” a mismatch of $\bm{U}(\Upsilon_{I},\psi_{mmI})$, while the second cavity sees a mismatch of $\bm{U}(\Upsilon_{12},\psi_{mm12})\bm{U}(\Upsilon_{I},\psi_{mmI})$. This means that, outside of the resonance of the second cavity ($r_{2}\simeq 1$), the response $\bm{F}_{2FC}$ is independent of $\Upsilon_{12}$ and its upper left element can be approximated by Eq. 8. Thus at the resonance of the first cavity, we expect similar degradation than the CFC case (QNR of 5 dB for the values in Table 1). On the contrary, outside of the resonance of the first cavity ($r_{1}\simeq 1$), $F_{2FC}$ can be approximated by $F_{1FC}$, but provided that one makes the substitution $\bm{U}(\Upsilon_{I},\psi_{mmI})\to\bm{U}(\Upsilon_{12},\psi_{mm12})\bm{U}(\Upsilon_{I},\psi_{mmI})$. We then redefine $\alpha=1-2\Upsilon^{\prime}+2\beta^{\prime}\sqrt{\Upsilon^{\prime}\Upsilon_{O}}e^{\mathrm{i}\psi^{\prime}}$ where $\beta^{\prime}=\sqrt{\frac{1-\Upsilon^{\prime}}{1-\Upsilon_{O}}}$ and the expressions of $\Upsilon^{\prime}$ and $\psi^{\prime}$ are given in Eq. 66 of Appendix E. Importantly, the maximum amount of mismatching thus obtained is larger than the sum of the mismatches: $\Upsilon^{\prime}_{\text{max}}=0.088>\Upsilon_{I}+\Upsilon_{12}=0.05$, which means that the degradation at the second cavity’s resonance can be larger than the first’s, as shown in the figure. Numerically, at $\Delta\omega_{2}$, we find, by replacing $\alpha$ with $\alpha^{\prime}$ in the formulas for 1FC, that in the worst case ($\Upsilon^{\prime}=\Upsilon^{\prime}_{\text{max}}$ and $\psi^{\prime}=\pi$), $\eta[\Delta\omega_{2}]=0.74$ and $\Xi[\Delta\omega_{2}]=0.025$, implying that QNR $\sim 3$ dB, as shown in Fig. 5. Importantly, this highlights that mode mismatch does not affect both cavities similarly, as the second cavity may experience larger degradation around its frequencies of resonance. This means that, in order to favor low frequencies, it may be beneficial to place the cavity with the lowest detuning first, opposite to what is shown in Fig. 5.
III.3 Filter cavity length fluctuations
The locking imprecision on the filter cavities induce a length fluctuation $\delta L$ which converts into phase noise for the squeezed field. Using the model developed in Section A.4, we plot its effect on squeezing in Fig. 6. The input squeezing level is 10 dB. We see that overall, cavity length fluctuations of less than 2 pm do not significantly degrade squeezing, neither for CFC nor 2FC, with a marginally better performance of CFC. Importantly, we have assumed that all the cavity lengths ($L_{1}$ and $L_{2}$ for the 2FC; $L_{c}$ and $L_{a}$ for the CFC) have the same fluctuation RMS $\delta L$. In practice, this will depend on the actual control scheme used to stabilize each of these cavities. While we defer the concrete technical control scheme to future work, we note that the range of values chosen for this length noise is realistic, as an FDS cavity controlled down to $\delta L\sim 1$ ppm was achieved in the current generation of GW detectors [9, 8].
Figure 6: Squeezing degradation for length fluctuation $\delta L$ for the 2FC and CFC, for a total length $L_{1}+L_{2}=L_{a}+L_{c}=5\text{km}+5\text{km}=10$ km.
III.4 Full degradation budget comparison
Figure 7: Comparison of CFC and 2FC using the full budget of Table 1.
Having individually studied all three loss contributions, we now determine the full picture of loss degradation in a 2FC and CFC. The response is illustrated in Fig. 7. The non-linear addition of mode mismatches in 2FC, along with the extra optical losses of the FI, ultimately degrade the total squeezing level below that of the CFC scheme. This study thus concludes that, all other parameters and degradation sources being equal, CFC overall performs better than 2FC.
IV Tuned single filter cavity configuration for Einstein Telescope
It has been repeatedly noted in the literature that the practical operation of a GW interferometer in a detuned configuration poses significant technical challenges in the stabilization of the various control loops [36, 37]. In this section, we propose an alternative to ET-LF that corresponds to a tuned configuration, along with a single filter cavity for optimal frequency-dependent squeezing (referred henceforth as the 1FC configuration). This would serve as either 1) an intermediary and easier to control configuration before upgrading to the full detuned setpoint or 2) a fallback solution, should detuned ET-LF prove to be technically difficult. Importantly, we emphasize that the 1FC configuration uses essentially the same infrastructure and excavation footprint as CFC, so that no significant overhead may be required, other than changing the mirror transmissivities. The optical layout is shown in Fig. 8 and compared to CFC: the length of the 1FC cavity is $L_{\text{FC}}=2L=L_{a}+L_{c}$.
Figure 8: Comparison of CFC (top) and 1FC (bottom) schemes, respectively for detuned and tuned ET-LF. One swaps from CFC to 1FC by removing the middle mirror and changing the input mirror transmissivity $T_{c}\to T_{\text{FC}}$.
Previous studies of tuned ET-LF with 1FC, such as [21, 38], only considered quantum noise degradation due to round trip loss and did not assume equal footprint $L_{\text{1FC}}=L_{a}+L_{c}$ in their comparison to detuned ET-LF + CFC. In the following, we provide a comparison with the full noise budget, which also includes contributions from the interferometer, and we keep the total 1FC/CFC lengths to be equal.
The setpoint and the noise curves for CFC are obtained using the method described in Section III, then converted to redshift units. To maximize the astrophysical reach of tuned ET + 1FC, we optimize on the redshift and use as free parameters the signal recycling transmissivity $T_{\text{SRM}}$, the transmissivity of the 1FC’s input mirror $T_{\text{FC}}$ as well as the detuning of 1FC $\Delta\omega_{\text{FC}}$. These are natural parameters to optimize [19, 39, 36, 28]. The results for total length of $L_{\text{1FC}}=L_{a}+L_{c}=2\text{ km}$ and $10$ km are presented in Fig. 9, while the optimized parameters for 1FC are shown in Table 2. Note that, thanks to the fact that the 1FC experiences less overall quantum degradation, it is possible to push the injected squeezing up to 19 dB without suffering from significant dephasing or phase noise.
Figure 9: Astrophysical reach of ET when ET-LF is tuned or not detuned, assuming the total length of 2 FC and length of 1FC will be identical. SEC refers to ”Signal Extraction Cavity”.
Parameter Physical meaning Value
$T_{\mathrm{SRM}}$ Transmissivity of signal recycling mirror $50$
$T_{\text{FC}}^{\mathrm{2km}}$ 2km FC input mirror transmissivity $0.08$
$\Delta\omega_{\text{FC}}^{\mathrm{2km}}$ 2km FC detuning $4.78$ Hz
$T_{\text{FC}}^{\mathrm{5km}}$ 10km FC input mirror transmissivity $0.40$
$\Delta\omega_{\text{FC}}^{\mathrm{10km}}$ 10km FC detuning $4.78$ Hz
$r_{\text{tuned}}$ Injected squeezing $19$ dB
Table 2: Optimized parameters for tuned ET-LF + 1FC.
[…]
As a conclusion, this study reinforces the relevance of aiming for the tuned + 1FC configuration as an alternative or an intermediary for the detuned configuration.
V Conclusions and future steps
Frequency dependent squeezing has been proved an effective techniques to reduce quantum noise in gravitational-wave detectors. In this work we have extended the previous studies about FDS requiring complex cavity systems to achieve the full rotation of the squeezing ellipse, and applied the results to ET-LF. We have seen that, while the constraints on the CFC parameters are stringent, they are readily achievable and imperfections can be partially compensated. Importantly, the CFC configuration behaves equally or better in terms of robustness with respect to squeezing degradation sources compared to the 2FC one. In particular, we have provided new theoretical insights in understanding cascaded mode mismatch on the degradation of squeezed light. For future work, special attention should nevertheless be given to the realization of the middle mirror’s parameters, and controllability issues of the CFC scheme should be addressed for this precise configuration, expanding for example on [40]. We also note that our insights on quantum degradation in multiple cavity schemes could be extended to hybrid EPR+1FC [22] or quantum teleportation schemes [20], where the interferometer itself acts as an effective single or multi-cavity.
Armed with the detailed model of quantum noise, we have studied the tuned ET-LF + 1FC configuration – similar to the layout of LIGO during the O4 run – and compared it to the final detuned ET-LF + FDS design. This comparison is relevant because the footprint of both layouts are similar, a crucial point for an underground interferometer. The relatively modest redshift gain of the detuned configuration compared to the tuned case incites us to consider a 2-step implementation for ET: first attain the tuned configuration by removing the middle mirror of CFC, then, add the latter and detune the interferometer to achieve the designed ET-LF sensitivity. Such an approach, should its controllability be demonstrated, may simplify the overall optical scheme, and mitigate the effects of some squeezing degradation sources.
Acknowledgements
We thank Michel Lequime, Myriam Zerrad, Anne Daumas, Michael Hartmann, Giancarlo Cella, Jérôme Degallaix, and Matthew Evans for the fruitful discussions and comments. This work has been supported by ANR-23-CE31-0004, by the Paris Region and by the EU Horizon 2020 Research and Innovation Program under the Marie Sklodowska-Curie Grant Agreement No. 101003460 ”PROBES”.
Appendix A Formalism
In this appendix, we describe the mathematical formalism used in the main sections of the paper.
A.1 Notations, Conventions
A classical, monochromatic electric field at frequency $\omega>0$, which propagates along $(+x)$, is written in complex form $E(t,x)=E_{0}\,\exp\left(-i\omega t+ikx\right)$ with $k=\frac{\omega}{c}$. In the presence of a cavity, we decompose the optical frequency $\omega=\omega_{0}+\Omega=n\omega_{\text{FSR}}+\Delta\omega+\Omega$ where $\omega_{0}$ is the reference frequency of the bright coherent field (1550nm for ET LF), $\omega_{\text{FSR}}$ is the free spectral range of the cavity, $\Delta\omega$ the detuning of the cavity, $\Omega$ the sideband frequency with respect to $\omega$, and $n\in\mathbb{N}$.
A.2 Matrix formalism in passive systems
Let us consider two copropagating spatial modes $a_{1}$ and $a_{2}$ (for example TEM00 and TEM01) of the electromagnetic field, written in the ladder basis vector $\bm{a}[\Omega]=\begin{bmatrix}a_{1}[\Omega]\\ a_{2}[\Omega]\end{bmatrix}$. Then, the usual transformations are
  1. 1.
    Free propagation without loss:
    $\bm{L}(\Omega,\Delta L,\psi)=\exp\left(i\frac{\omega_{0}\Delta L+\Omega L_{0}}{c}\right)\begin{bmatrix}1&0\\ 0&e^{i\psi}\end{bmatrix}$ (18)
    where $\Delta L$ is the microscopic length detuning from the resonance, $L_{0}$ is the macroscopic length of the cavity on resonance, $\psi$ is the Gouy phase of the higher order mode $a_{2}$ relative to $a_{1}$.
  2. 2.
    Power loss:
[…]
    where $\Lambda$ is the power loss.
  3. 3.
    Mirror reflection and transmission:
    $\bm{r}=\sqrt{1-T}\,\bm{1}\qquad\text{and}\qquad\bm{t}=\sqrt{T}\,\bm{1}$ (20)
    where $T$ is the power transmissivity.
  4. 4.
    Mode mismatch:
    $\bm{U}(\Upsilon,\psi_{mm})=\begin{bmatrix}\sqrt{1-\Upsilon}&-\sqrt{\Upsilon}e^{\mathrm{i}\psi_{mm}}\\ \sqrt{\Upsilon}e^{-\mathrm{i}\psi_{mm}}&\sqrt{1-\Upsilon}\\ \end{bmatrix}$ (21)
    where $\Upsilon$ is the amplitude of the mismatch, while $\psi_{mm}$ is the phase of the mismatch.
A.3 Quantum noise in passive systems
Consider a system which is only composed of passive elements. Its classical transfer function on the signal mode is
$a^{\text{out}}[\Omega]=F[\Omega]a^{\text{in}}[\Omega]$ (22)
so in the two photon formalism, its matrix has the form
$\bm{\mathfrak{a}}^{\text{out}}[\Omega]=\begin{bmatrix}F[\Omega]&0\\ 0&F[-\Omega]^{*}\end{bmatrix}\bm{\mathfrak{a}}^{\text{in}}[\Omega]=\bm{F}[\Omega]\bm{\mathfrak{a}}^{\text{in}}[\Omega]$ (23)
We recall that we can go from the sideband basis to the quadrature basis using the isomorphism:
$\bm{x}^{\text{out}}=\bm{G}\bm{x}^{\text{in}}=\bm{PFP^{\dagger}}\bm{x}^{\text{in}}$ (24)
where $\bm{P}=\frac{1}{\sqrt{2}}\begin{bmatrix}1&1\\ -i&i\end{bmatrix}$.
Finally, to make the spectral density matrix (SDM) compatible with the added quantum noise, the transformation is [33](we use single sided power spectral densities)
$\bm{\bar{S}}^{\text{out}}=\bm{G}\left(\bm{\bar{S}}^{\text{in}}-\bm{1}\right)\bm{G}^{\dagger}+\bm{1}$ (25)
Analytically, $\bm{G}[\Omega]$ has the expression
$\bm{G}[\Omega]=e^{\mathrm{i}\theta_{C}}\left(C[\Omega]\bm{1}-D[\Omega]\bm{\sigma}\right)\bm{R}(\theta_{D})$ (26)
where
$C[\Omega]$ $=\frac{|F[\Omega]|+|F[-\Omega]|}{2}$ (27)
$D[\Omega]$ $=\frac{|F[\Omega]|-|F[-\Omega]|}{2}$ (28)
$\theta_{C}[\Omega]$ $=\frac{\arg(F[\Omega])-\arg(F[-\Omega])}{2}$ (29)
$\theta_{D}[\Omega]$ $=\frac{\arg(F[\Omega])+\arg(F[-\Omega])}{2}$ (30)
$\bm{\sigma}$ $=\begin{bmatrix}0&-\mathrm{i}\\ \mathrm{i}&0\end{bmatrix}$ (31)
We assume that the input SDM is a squeezed rotated state:
$\bm{\bar{S}}^{\text{in}}=\bm{R}(\phi)\bm{S}(2r)\bm{R}(\phi)^{T}$ (32)
with $\bm{S}(r)=\text{diag}(e^{r},e^{-r})$ the squeezing matrix. Noting that $\bm{R}(\theta_{D})$ commutes with $\bm{\sigma}$ and that the latter matrix is hermitian, we find that
$\bm{\bar{S}}^{\text{out}}=\bm{R}(\theta_{D}+\phi)\left[(C\bm{1}-D\bm{\sigma})(\bm{S}(2r)-\bm{1})(C\bm{1}-D\bm{\sigma})+\bm{1}\right]\bm{R}(\theta_{D}+\phi)^{T}$ (33)
The properties of this matrix are better understood by factoring out the rotation matrix, that is, define $\bm{\bar{S}^{\prime}}\coloneqq\bm{R}(\theta_{D}+\phi)^{T}\bm{\bar{S}}^{\text{out}}\bm{R}(\theta_{D}+\phi)$ so that
$\bm{\bar{S}^{\prime}}=(C\bm{1}-D\bm{\sigma})(\bm{S}(2r)-\bm{1})(C\bm{1}-D\bm{\sigma})+\bm{1}$ (34)
Explicitly, we find
$\bm{\bar{S}^{\prime}}=\begin{bmatrix}\eta\left[(1-\Xi)e^{2r}+\Xi e^{-2r}\right]+(1-\eta)&4\mathrm{i}CD\sinh^{2}(r)\\ -4\mathrm{i}CD\sinh^{2}(r)&\eta\left[\Xi e^{2r}+(1-\Xi)e^{-2r}\right]+(1-\eta)\end{bmatrix}$ (35)
where
$\eta$ $\coloneqq C^{2}+D^{2}=\frac{|F[\Omega]|^{2}+|F[-\Omega]|^{2}}{2}$ (36)
$\Xi$ $\coloneqq\frac{D^{2}}{\eta}=\frac{1}{2}-\frac{|F[\Omega]F[-\Omega]|}{|F[\Omega]|^{2}+|F[-\Omega]|^{2}}$ (37)
Thus the squeezing quadratures are rotated by an amount equal to the frequency-dependent angle $\theta_{D}[\Omega]$. Degradation mechanisms may affect this angle; we thus define misphasing $\Delta\theta_{D}[\Omega]$ as the difference between $\theta_{D}[\Omega]$ when there is no loss in the passive system and the lossy case:
$\Delta\theta_{D}[\Omega]\coloneqq\theta_{D}[\Omega]\big{|}_{\text{lossless}}-\theta_{D}[\Omega]$ (38)
The expression Eq. 35 helps one understand how squeezing is degraded when doing homodyne detection. The general expression of the homodyne photocurrent spectrum is [33]
$\bar{S}_{\theta}^{\text{HD}}[\Omega]=\bm{u}(\theta)^{T}\text{Re}(\bm{\bar{S}^{\prime}}[\Omega])\bm{u}(\theta)=\Lambda_{+}^{\mathbb{R}}\cos^{2}(\theta)+\Lambda_{-}^{\mathbb{R}}\sin^{2}(\theta)$ (39)
where the homodyne readout is $u(\theta)=\begin{bmatrix}\cos(\theta)\\ \sin(\theta)\end{bmatrix}$, and $\Lambda_{\pm}^{\mathbb{R}}\coloneqq\eta\left[(1-\Xi)e^{\pm 2r}+\Xi e^{\mp 2r}\right]+(1-\eta)$ are the eigenvalues of $\text{Re}(\bm{\bar{S}^{\prime}})$. This can be seen by noting that it is a diagonal matrix per Eq. 35.
The interferometer is designed to rotate the squeezed field by amount $\pi/2-\theta_{D}[\Omega]\big{|}_{\text{lossless}}$. Effectively, this means that the equivalent homodyne angle is $\theta=\Delta\theta_{D}[\Omega]+\pi/2$, so the measured squeezing through homodyne detection is
$\bar{S}^{\text{HD}}[\Omega]=\Lambda_{+}^{\mathbb{R}}\sin^{2}(\Delta\theta_{D}[\Omega])+\Lambda_{-}^{\mathbb{R}}\cos^{2}(\Delta\theta_{D}[\Omega])$ (40)
which is exactly Eq. 7.
A.4 Phase noise
Phase noise on a parameter $X_{k}$ on which the spectral density depends is accounted by effectively by averaging its expression in the interval $[X_{k}-\delta X_{k},X_{k}+\delta X_{k}]$:
$\begin{split}\bar{S}^{\text{avg}}=&\bar{S}(X_{1},\dots,X_{n})\\ &+\sum_{k}\left(\frac{\bar{S}(X_{k}+\delta X_{k})+\bar{S}(X_{k}-\delta X_{k})}{2}-\bar{S}(X_{k})\right)\\ \end{split}$ (41)
where in the sum, the dependence of $\bar{S}$ on all the other parameters $X_{j}$ for $j\neq k$ are made implicit and evaluated at the setpoint [41].
Appendix B Proof of the scaling laws for bandwidth, detuning and loss
In this appendix we prove the equations Eqs. 1, 2, 4 and 3 as well as the observation of Section III.1, that the responses of 2FC and CFC are similar when their RTL $\Lambda$ and $\Lambda^{\prime}$ are equal. Our proof generalizes that of [22] so this latter observation.
B.1 Single filter cavity
We recall that the frequency response of a single filter cavity
$r_{\text{cav}}=\frac{r-\sqrt{1-\Lambda}\,\mathrm{e}^{-2\mathrm{i}\varphi}}{1-{r}\sqrt{1-\Lambda}\,\mathrm{e}^{-2\mathrm{i}\varphi}}$ (42)
where the input mirror amplitude reflectivity is $r$, the RTL is $\Lambda$, the single-trip phase is $\varphi\coloneqq(\Delta\omega+\Omega)L_{0}/c=(\omega_{0}\Delta L+\Omega L_{0})/c$. Expanding at lowest order in the input mirror transmission $T$, $\Lambda$ and $\phi\coloneqq-2\varphi$ gives $r=\sqrt{1-T}\simeq 1-\frac{T}{2}$, $\sqrt{1-\Lambda}\simeq 1-\frac{\Lambda}{2}$, $\mathrm{e}^{-2\mathrm{i}\varphi}\simeq 1+\mathrm{i}\phi$. Hence
$r_{\text{cav}}\simeq\frac{-T+\Lambda-2\,\mathrm{i}\phi}{T+\Lambda-2\,\mathrm{i}\phi}$ (43)
B.2 Two cavities
The 2FC configuration is simply the product of the responses of two single cavities (indices 1 and 2). Assuming that the RTL $\Lambda\ll T_{1,2}$ correspond to mirror absorption loss, and that they are the same for all mirrors in the setup,
$r_{\text{2FC}}$ $=\frac{-T_{1}+\Lambda-2\,\mathrm{i}\phi_{1}}{T_{1}+\Lambda-2\,\mathrm{i}\phi_{1}}\times\frac{-T_{2}+\Lambda-2\,\mathrm{i}\phi_{2}}{T_{2}+\Lambda-2\,\mathrm{i}\phi_{2}}$ (44)
$=\frac{\frac{T_{1}T_{2}}{4}+\frac{\mathrm{i}}{2}[\phi_{1}(T_{2}-\Lambda)+\phi_{2}(T_{1}-\Lambda)]-\phi_{1}\phi_{2}}{\frac{T_{1}T_{2}}{4}-\frac{\mathrm{i}}{2}[\phi_{1}(T_{2}+\Lambda)+\phi_{2}(T_{1}+\Lambda)]-\phi_{1}\phi_{2}}$ (45)
B.3 Coupled Filter Cavity with symmetric loss
For a coupled filter cavity, the response is
$r_{\text{CFC}}=\frac{r_{c}-r_{\text{ref}}\,\sqrt{1-\Lambda^{\prime}}\,\mathrm{e}^{-2\mathrm{i}\varphi_{c}}}{1-r_{\text{ref}}\,r_{c}\sqrt{1-\Lambda^{\prime}}\,\mathrm{e}^{-2\mathrm{i}\varphi_{c}}}$ (46)
where $r_{\text{ref}}$ is the response of the a-cavity, and we have written $\Lambda^{\prime}$ the RTL. It is a priori different from $\Lambda$ but we assume it to be identical for both sub-cavities of the CFC system. Thus the reflectivity of the a-cavity is
$r_{\text{ref}}=\frac{r_{a}-\sqrt{1-\Lambda^{\prime}}\,\mathrm{e}^{-2\mathrm{i}\varphi_{a}}}{1-{r_{a}}\sqrt{1-\Lambda^{\prime}}\,\mathrm{e}^{-2\mathrm{i}\varphi_{a}}}$ (47)
We inject in the previous equation to obtain
$r_{\text{CFC}}=\frac{r_{c}-r_{a}\sqrt{1-\Lambda^{\prime}}\,(r_{c}\mathrm{e}^{-2\mathrm{i}\varphi_{a}}+\mathrm{e}^{-2\mathrm{i}\varphi_{c}})+(1-\Lambda^{\prime})\,\mathrm{e}^{-2\mathrm{i}\varphi_{a}}\,\mathrm{e}^{-2\mathrm{i}\varphi_{c}}}{1-r_{a}\sqrt{1-\Lambda^{\prime}}\,(\mathrm{e}^{-2\mathrm{i}\varphi_{a}}+r_{c}\,\mathrm{e}^{-2\mathrm{i}\varphi_{c}})+(1-\Lambda^{\prime})r_{c}\,\mathrm{e}^{-2\mathrm{i}\varphi_{a}}\,\mathrm{e}^{-2\mathrm{i}\varphi_{c}}}$ (48)
Using the lowest-order approximations, the numerator becomes, after simplification
$\text{Num}(r_{\text{CFC}})=T_{a}+\frac{\mathrm{i}}{2}[(T_{c}+T_{a}-\Lambda^{\prime})\phi_{a}+(T_{a}-\Lambda^{\prime})\phi_{c}]-\phi_{a}\phi_{c}$ (49)
Comparing this with the numerator of Eq. 45, and identifying the real and imaginary parts, we get
$T_{a}-\phi_{a}\phi_{c}$ $=\frac{T_{1}T_{2}}{4}-{\phi_{1}\phi_{2}}$ (50)
$(T_{c}+T_{a}-\Lambda^{\prime})\phi_{a}+(T_{a}-\Lambda^{\prime})\phi_{c}$ $=\phi_{1}(T_{2}-\Lambda)+\phi_{2}(T_{1}-\Lambda)$ (51)
We now expand the expressions of $\phi_{i}=-2(\Omega+\Delta\omega_{i})\tau$ for $i\in\{1,2,a,c\}$, while $\tau=L/c$ where $L$ is the length the cavity, assumed to be identical for all four cavities. We then identify both sides of Eqs. 50 and 51 by powers of $\Omega$ and assuming $T_{a}\ll T_{c}$:
$\gamma_{a}$ $=\tau_{c}(\gamma_{1}\gamma_{2}-\Delta\omega_{1}\Delta\omega_{2}+\Delta\omega_{a}\Delta\omega_{c})$ (52)
$\Delta\omega_{a}+\Delta\omega_{c}$ $=\Delta\omega_{1}+\Delta\omega_{2}$ (53)
$2(\gamma_{c}+\gamma_{a}-\gamma_{2}-\gamma_{1})$ $=\frac{\Lambda^{\prime}-\Lambda}{\tau}$ (54)
$4(\gamma_{c}\Delta\omega_{a}+\gamma_{a}\Delta\omega_{c}-\gamma_{2}\Delta\omega_{1}-\gamma_{1}\Delta\omega_{2})$ $=\frac{(\Delta\omega_{1}+\Delta\omega_{2})(\Lambda^{\prime}-\Lambda)}{\tau}$ (55)
where the bandwidths $\gamma_{i}=\frac{cT_{i}}{4L}$ correspond to the ones in the lossless case. Thus, the scaling laws Eqs. 1, 2, 4 and 3 are obtained by solving the previous equations assuming $\Lambda^{\prime}=0$ and $\Lambda=0$. Then, we obtain in the lossy case that $\Lambda^{\prime}=\Lambda$ by looking at Eqs. 54 and 55, the left sides being zero. This concludes the proof of Eqs. 1, 2, 4 and 3 and formally justifies the observation of Section III.1.
Appendix C Substrate loss of the middle mirror in CFC
In this appendix we justify the statement of Section III.1 that substrate loss of the middle mirror in CFC has a negligible impact on the degradation of squeezing. We consider a middle mirror made of fused silica and 5 cm thick. Substrate loss in Fused Silica core fibers are of the order of 1 ppm / cm [42, 43]. We take a conservative 2 ppm/cm figure for bulk fused silica, which amounts to an added round trip loss of 20 ppm. Using this value, and adding the 30 ppm of loss per subcavity (see Table 1), we compare the squeezing performances when the substrate is facing towards the c-cavity (with High Reflectivity coating faced towards the a-cavity) to when it is facing the a-cavity. From the results in Fig. 10, we see that by placing the substrate in the c-cavity, squeezing is never degraded by more than 0.5 dB compared to the 2FC scheme.
Figure 10: Comparison of CFC vs 2FC when only optical losses are considered, with extra 20 ppm round trip loss in CFC due to the substrate of the middle mirror.
Appendix D Mode mismatch in a coupled filter cavity
In this appendix we compute the mode mismatch between the two sub-cavities of a CFC setup. We consider the configuration relevant to this paper, that is, in the perfect case, both input and end mirrors (indices $c$ and $e$) are convex with the same radius of curvature $R_{a}=R_{c}=R$, are separated by a length $L$ such that $L<2R$ (for stability reasons) and the middle mirror (index $a$)is flat at the position $L/2$. This forms two plano-concave cavities, with identical waist position (at the position of the middle mirror) and waist size (since the input and end mirrors have the same RoC).
Now, for a realistic cavity, assume that the radii of curvatures are $R_{c}=R(1+\epsilon_{c})$, $\infty>R_{a}\gg R$ and $R_{e}=R(1+\epsilon_{e})$, with $\epsilon_{a,e}\ll 1$. The middle mirror is still placed at $L/2$.
Since we are interested in the relative mode mismatch between the two subcavities, we can assume without loss of generality that $\epsilon_{c}=0$ and concentrate all the defects on the parameter $\epsilon_{e}$ of the end mirror. This is represented in Fig. 11.
Figure 11: Realistic CFC layout with mode mismatch. The upper figure shows the CFC with perfect mirrors: the waist of the two subcavities is at the position of the middle mirror., and both curved mirrors have the same radius of curvature. The bottom figure shows imperfections in the a-cavity.
Under this assumption, the fundamental eigenmode of the first sub-cavity is the same as in the perfect case, while the second cavity mode is changed. To determine the mode mismatch (i.e. the mode overlap between the two fundamental eigenmodes), we first compute the position and the size of the waist of the second sub-cavity.
First, let $z=0$ be the middle of the cavity, at the position of the middle mirror. The waist position of the second cavity, $z^{\prime}_{0}$, is given by
$z^{\prime}_{0}=\frac{d(R(1+\epsilon_{e})-d)}{R_{a}+R(1+\epsilon_{e})-2d}$ (56)
where a positive RoC value corresponds to a concave mirror and $d=L/2$. The waist squared is
$w_{0}^{\prime 2}=\frac{\lambda}{\pi}\sqrt{z_{0}^{\prime}(R_{a}-z_{0}^{\prime})}$ (57)
In the perfect case, the waist position and size are
$z_{0}=0,\quad w_{0}^{2}=\frac{\lambda}{\pi}\sqrt{d(R-d)}$ (58)
In the imperfect case, at lowest order in $\epsilon$ and $\frac{1}{R_{a}}$, the waist position is
$z^{\prime}_{0}=\frac{d(R-d)}{R_{a}}$ (59)
and the waist size is
$w^{\prime}_{0}=w_{0}\left(1+\epsilon\frac{R}{4(R-d)}-\frac{R-2d}{4R_{a}}\right)$ (60)
The complex mismatch $m=\sqrt{\Upsilon_{c}}\,e^{\mathrm{i}\psi_{mmc}}$ of the cavity is [35]:
$m$ $=\frac{w^{\prime}_{0}-w_{0}}{w_{0}}+\mathrm{i}\frac{z_{0}^{\prime}}{2\sqrt{d(R-d)}}$ (61)
$=\epsilon\frac{R}{4(R-d)}-\frac{R-2d}{4R_{a}}+\mathrm{i}\frac{\sqrt{d(R-d)}}{2R_{a}}$ (62)
from which we identify $\Upsilon_{c}=|m|^{2}$ and $\psi_{mmc}=\arg(m)$.
Numerically, state of the art mirrors fabrication allows to obtain for a flat mirror $R_{a}>10^{9}$ m. This lower bound is estimated by making the assumption that the RMS flatness $\sigma\sim 10^{-10}$ m of a state-of-the art mirror is due to its sphericity. Then the radius of curvature is of the order of $D^{2}/\sigma$ where $D\sim 0.3$ m is the diameter of the mirror. In practice, since mirror maps show no quadratic contribution to $\sigma$ (see [24]), we have $R_{a}>D^{2}/\sigma$. We further have $R\epsilon\sim 1$ m, for typical $R=15$ km. We thus expect $\Upsilon_{c}\sim 10^{-9}$ for $L=1-10$ km. This shows that, provided state-of-the art mirror quality, the internal mode mismatch of a coupled cavity is negligible.
Likewise, if the length of the second cavity changes from $d$ to $d+\delta$ with $\delta\ll d$, the effect at lowest order is to change the waist size to be
$w_{0}^{\prime}=w_{0}+\left(1+\frac{\delta R}{4d(R-d)}\right)$ (63)
which means that the mode mismatch gets the contribution
$m=\frac{w^{\prime}_{0}-w_{0}}{w_{0}}=\frac{\delta R}{4d(R-d)}$ (64)
For $\delta\sim 1$ m and the same values for $R$ and $d$ as previously, we get $m\sim 10^{-4}$ so $\Upsilon_{a}\sim 10^{-8}$. Again, this shows that the mismatch due to cavity length defects is negligible.
Appendix E Addition of mode mismatches
Let $\Upsilon_{A},\psi_{A}$ and $\Upsilon_{B},\psi_{B}$ be mismatch parameters. Let $\bm{U}_{C}\coloneqq\bm{U}(\Upsilon_{A},\psi_{A})\bm{U}(\Upsilon_{B},\psi_{B})$. We wish to write $\bm{U}_{C}$ in a form similar to $\bm{U}(\Upsilon^{\prime},\psi^{\prime})$ where $(\Upsilon^{\prime},\psi^{\prime})$ can be expressed as a function of $\Upsilon_{A},\psi_{A},\Upsilon_{B},\psi_{B}$.
The product yields
$\bm{U}_{C}=\begin{bmatrix}\sqrt{1-\Upsilon^{\prime}}\,e^{\mathrm{i}\psi_{0}}&-\sqrt{\Upsilon^{\prime}}\,e^{\mathrm{i}\psi_{1}}\\ \sqrt{\Upsilon^{\prime}}\,e^{-\mathrm{i}\psi_{1}}&\sqrt{1-\Upsilon^{\prime}}\,e^{-\mathrm{i}\psi_{0}}\end{bmatrix}$ (65)
where
$\begin{split}\Upsilon^{\prime}=&\Upsilon_{A}+\Upsilon_{B}-2\Upsilon_{A}\Upsilon_{B}\\ &+2\sqrt{\Upsilon_{A}(1-\Upsilon_{B})\Upsilon_{B}(1-\Upsilon_{A})}\,\cos(\psi_{A}-\psi_{B})\end{split}$ (66)
$\psi_{0}=\arg(\sqrt{(1-\Upsilon_{A})(1-\Upsilon_{B})}-\sqrt{\Upsilon_{A}\Upsilon_{B}}\,e^{\mathrm{i}(\psi_{A}-\psi_{B})})$ and $\psi_{1}=\arg(\sqrt{(1-\Upsilon_{A})\Upsilon_{B}}\,e^{\mathrm{i}\psi_{B}}+\sqrt{(1-\Upsilon_{B})\Upsilon_{A}}\,e^{\mathrm{i}\psi_{A}})$. Defining $\psi^{\prime}\coloneqq\psi_{1}-\psi_{0}$, we get that
$\bm{U}_{C}=e^{\mathrm{i}\psi_{0}}\,\bm{U}(\Upsilon^{\prime},\psi^{\prime})\,\text{diag}(1,e^{-2\mathrm{i}\psi_{0}})$ (67)
Note that on the right hand side, the complex exponential adds a global phase so is irrelevant, while the diagonal matrix’s non trivial bottom right component is usually applied to the higher order mode in a vacuum state, whose properties do not depend on the phase either, so can also be neglected. Under these assumptions, we thus have $\bm{U}_{C}\sim\bm{U}(\Upsilon^{\prime},\psi^{\prime})$.
We see that $\Upsilon^{\prime}$ is maximized when $\psi_{A}=\psi_{B}$ (constructive interference of the higher order mode) and equals
$\Upsilon^{\prime}_{\text{max}}=\left(\sqrt{\Upsilon_{A}(1-\Upsilon_{B})}+\sqrt{\Upsilon_{B}(1-\Upsilon_{A})}\right)^{2}$ (68)
while its minimum is attained for $\psi_{A}=\psi_{B}+\pi$ and equals
$\Upsilon^{\prime}_{\text{min}}=\left(\sqrt{\Upsilon_{A}(1-\Upsilon_{B})}-\sqrt{\Upsilon_{B}(1-\Upsilon_{A})}\right)^{2}$ (69)
These expressions can be simplified by defining the polar variables $\theta_{A}$ and $\theta_{B}$ such that $\sin(\theta_{A})=\sqrt{\Upsilon_{A}}$, $\cos(\theta_{A})=\sqrt{1-\Upsilon_{A}}$ and similarly for $\theta_{B}$. Then,
$\Upsilon^{\prime}_{\text{max}}=\sin^{2}(\theta_{A}+\theta_{B})\quad\text{and}\quad\Upsilon^{\prime}_{\text{max}}=\sin^{2}(\theta_{A}-\theta_{B})$ (70)